OpenStudy (anonymous):

What is the similarity ratio of the smaller to larger similar cylinders? smaller cylinder r=12 height=20 larger cylinder r=18 height=30

OpenStudy (mathstudent55):

What is \(\dfrac{12}{18} = ~? \) What is \(\dfrac{20}{30} = ~? \) Reduce the fractions. they will both give you the same result.

OpenStudy (anonymous):

12/18=.66 20/30=.66

OpenStudy (anonymous):

so is that the ratio?

OpenStudy (jhannybean):

Yep. or you can put it in fraction form, \(\large \frac{2}{3}\)